OpenStudy (anonymous):

y=mx+b Slope 3/2, passing through (-2,0) What is the y intersept?

OpenStudy (anonymous):

use the point slope form y-y1=m(x-x1) y-0=3/2(x-(-2)) y=(3/2)x+3 so 3 would be the y intercept

OpenStudy (anonymous):

Oh, okay. Thanks

OpenStudy (anonymous):

no prob